OpenStudy (anonymous):

Four times a number plus five more than three times the number is 47.

OpenStudy (anonymous):

let the number be x 4x+(5+3x)=47 therefore 4x+3x=47-5=42 therefore 7x=42 therefore x=42/7 hence finally the number is 6